• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:01:36 on June 3,2002

    Entry number 85145

    keyword=Shift Summary

    Day Shift June 3, 2002
    John LeRose (SL)
    Wendy Hinton (TO)
    Serge Kox (AM)
    Ron Ransome (PM)

    No beam at start of shift
    MCC working on resteering in one of the Arcs
    Hall C Target cooldown planned for this morning
    Ready to take data at Q1_d20 (started earlier)
    9:10 MCC announces intent to start trying to deliver high current CW
    9:32 MCC trying again to send CW down the Hall A line
    9:37 trying tune beam instead
    10:50 still no beam! MCC says they think they've found the problem in the injector and it will take about 30 minutes to fix it. Concurrently, they've developed some other hardware problems (rf down etc.) which they will be working on at the same time. They hope to be ready to try again to deliver beam in about 30 minutes.
    12:13 NOW they say they're ready to give us beam (95 microamps?) Got some not useful beam in fits and starts for about 8 minutes (up to 50 microamps).
    12:38 MCC trying again to deliver CW beam.
    13:00 Started run 1128 @45 microamps
    13:19 MCC still getting BCM (BLM?) trips as they try to run the current up. They're giving it back to the injector folks to work on it some more. Run 1128 ended
    13:35 MCC ready to try again
    14:00 Beam on and off at up to 65 microamps for the last 25 minutes (FB off x=-0.21 or larger) Requested FB on and x closer to zero. They say they're still optimizing.
    15:10 requested access to check target auxilliary heater (Kathy McCormick and Dave Meekins)
    15:40 Something goes "pop" in the center room while Kathy is tinkering there. Doesn't smell too good either!
    15:52 Access ended. Auxilliary heaters are no more! See related halog entry.
    15:55 Return to beam permit
    Not a good shift!
    -JJL